# Investigating potential causes of Sepsis with Bayesian network structure learning

### v1

@@ Abstract
Sepsis is a life-threatening and serious global health issue. This study combines knowledge with available hospital data to investigate the potential causes of Sepsis that can be affected by policy decisions. We investigate the underlying causal structure of this problem by combining clinical expertise with score-based, constraint-based, and hybrid structure learning algorithms. A novel approach to model averaging and knowledge-based constraints was implemented to arrive at a consensus structure for causal inference. The structure learning process highlighted the importance of exploring data-driven approaches alongside clinical expertise. This includes discovering unexpected, although reasonable, relationships from a clinical perspective. Hypothetical interventions on Chronic Obstructive Pulmonary Disease, Alcohol dependence, and Diabetes suggest that the presence of any of these risk factors in patients increases the likelihood of Sepsis. This finding, alongside measuring the effect of these risk factors on Sepsis, has potential policy implications. Recognising the importance of prediction in improving Sepsis related health outcomes, the model built is also assessed in its ability to predict Sepsis. The predictions generated by the consensus model were assessed for their accuracy, sensitivity, and specificity. These three indicators all had results around 70%, and the AUC was 80%, which means the causal structure of the model is reasonably accurate given that the models were trained on data available for commissioning purposes only.
@@ 1 Introduction
Extensive research has been produced in building predictive models for the early diagnosis of Sepsis. Some examples are [4], [5], [6], [7], [8] and [9]. These models, which include Bayesian Networks and some constructed using supervised learning models, are intended to aid the management of Sepsis in hospitals and prevent severe cases by supporting the early detection of Sepsis. Although these models may help to reduce the incidence and mortality of Sepsis due to their early prediction capabilities, they do not investigate potential areas for intervention to reduce the incidence of Sepsis. Furthermore, all the predictive models reviewed use bio-markers or clinical data for predicting Sepsis. However, these data are not routinely collected by the National Health Service (NHS) in England for service commissioning purposes. This lack of data may be true for commissioners of services in other healthcare systems. Accessing patient records without a clinical/direct care reason is generally restricted. Even if commissioners could access these data, they probably would not have the necessary expertise to interpret it if they did not have a clinical background. Because commissioners play a crucial role in service design, building a model using data they regularly access, like the one developed in this study, is important for exploring new services that can improve patient outcomes.
@@
The methods used for prediction in the public health sector are generally regression-based. Two examples of this type of modelling were reviewed: [8] and [9]. One problem with regression models is that they do not enable accurate modelling of the often-interdependent relationships of the risk factors for patients developing Sepsis. The relationships between these factors might be highly complex, and it is essential to model them as such.
@@
Regression models do not generally explore possible causal relationships. In this study, we explore the graphical structures that structure learning algorithms extract from data and investigate whether these findings could indicate causal relationships. However, these algorithms rely on assumptions which often do not hold in practice. Here, we attempt to reduce the resulting errors by incorporating expert knowledge and averaging models produced by different algorithms.
@@
Causal Bayesian Networks (CBNs) are represented by a Directed Acyclic Graph (DAG), which provides an intuitive and easy way to read the causal relationships between risk factors and model the impact of interventions available to policymakers. CBNs consist of nodes representing variables and arcs assumed to represent causal relationships. The structure learning process described in the previous paragraph determines the structure of these arcs and nodes. One of the relevant features of CBNs is that they enable us to model predictive and diagnostic inference and the effect of hypothetical interventions using the so-called do-calculus framework proposed by [10].
@@
  • Develop a CBN model using data routinely collected by the NHS in England for commissioning purposes.
  • Capture the highly complex interrelations amongst risk factors with a novel method that combines expert knowledge and the averaging of six structure learning algorithms.
  • Assess the ability of the model to predict the presence of Sepsis.
  • Identify which risk factors can be influenced by policy decisions and use algorithms to discover and estimate their causal effects on Sepsis.
@@ 2.1 Factors Associated with Sepsis
A range of risk factors is identified in the literature as being associated with the presence of Sepsis. The risk factors described below can be found in routinely collected data for service commissioning purposes by the NHS. The discussion did not include risk factors that cannot be found in this data. Table 1 lists these factors and studies identifying them as likely to be linked with Sepsis. Our study will consider all of these factors.
@@
Table 2 presents other variables we incorporated based on clinical expertise in this study. Clinical expertise was elicited through collaboration with Dr Jonathon Dean, a clinical expert based at the North East London Adult Critical Care Operational Delivery Network and the co-founder of the North East London Critical Care Transfer and Retrieval service. Dr Dean has spent the last five years working in anaesthetics and critical care in North East London, where Sepsis is an important and topical pathology.
@@ 3 Data pre-processing
This study will use the Secondary Uses Service (SUS+) database, which includes data for patients admitted to hospitals (inpatients), patients using accident and emergency services, and patients with outpatient appointments within the NHS in England. This study will look at admissions classified as an emergency; this means admissions to hospitals that were not planned. ICD (International Classification of Diseases) 10 codes is the 10th version of the global standard for diagnosis classification ([19]). These are used to classify diagnoses in the database and will be crucial to constructing the variables identified by the literature review and expert knowledge. OPCS codes identify the different interventions, operations, and procedures within the NHS in England, and we have used them to create relevant variables in this study. Variables were constructed with clinical guidance, but only non-clinical authors had access to the database. SUS+ was accessed through the National Commissioning Data Repository (NCDR), a pseudonymised patient-level data repository managed by NHS England.
@@ 3.1 Data Extraction
The odbc R package ([20]) was used to connect to the database. This package also allows for the database to be queried using SQL. Each row in the admitted patients table represents an episode during a hospital spell. A spell can be composed of many episodes. The episodes of a spell are likely to obtain similar information regarding diagnosis, procedures and patient information. To obtain only one episode representing any given spell, the rule of the most relevant episode during a spell was implemented for the data extraction, where the most relevant is given by the most relevant Healthcare Resource Groups (HRGs) during the spell. In this way, we ensure we are not double counting the same spells while getting the most important information about that spell. Treatments are grouped into the same HRGs when they are deemed to require similar levels of resources. Therefore, we selected the episode where the highest level of resources was used for treatment.
The data extracted is for the financial year 2019/2020, which ran from the 1st of April 2019 until the 31st of March 2020. Except for the variables representing the age, number of diagnoses, number of procedures, sex, and ethnicity of a patient, all the variables extracted were in the form of a binary indicator establishing the presence or absence of the risk factor. A sample of 1,000,000 episodes was taken due to server constraints; i.e., the runtime for this study on the server already spans multiple days with this sample size.
@@ 3.2 Missing Data
Structure learning algorithms generally require complete datasets to work. This is also the case with the algorithms employed in this study. However, the data collated contained missing values. It is worth noting that the pattern of missing data in our dataset is very small. The proportion of missing values in the variables age, ethnicity, and sex) is less than 0.6%, i.e. less than 6,000 rows of our 1,000,000 sample.
A common but often problematic solution to this problem is to delete the rows with missing values. This works well under the very strong assumption that the missing data are Missing Completely At Random (MCAR) and if sufficient sample size remains after deletion. We say that data are MCAR when missingness is unrelated to other variables and to observed and unobserved values of the variable itself. A much more likely assumption to make for missing data in this study is that data are Missing At Random (MAR). Data is MAR when the fact that data is missing is related to other variables in the dataset but not to the variable itself. The data we use in this study is a cross-section of hospital episodes. Therefore, it is more reasonable to assume that the missing data are MAR; i.e., if a missing value is present, it seems plausible that it will be related to other patient information. Because deleting rows with missing data is not a good option for this study, we performed data imputation using the missForest R package ([21], [22])]), which is suitable for both types of missingness, MCAR and MAR. This algorithm uses a Random Forest to impute missing values and does not require any parameters to be specified.
When missing data are related to a diagnosis or a procedure, the value 0 is manually imputed. This action was taken assuming that missing information on diagnosis or procedures is equivalent to no diagnosis or procedures. It is not possible to differentiate in the data when a clinician or coder was unable to record a diagnosis or when it was not recorded because it was not relevant to the episode. However, the course of action taken seems reasonable because if no information was available for a diagnosis, it probably was irrelevant to the reason behind the hospitalization. Therefore, although we use the language of missing data for these cases, there are strong arguments to not consider them missing but not applicable. Where the data records presented wrongly inputted values in the variables describing the age and ethnicity of a person, the following actions were taken:
  • For age, this was handled by converting the values over 120 to NA (indication of a missing value in R) and then using the imputation method chosen. This is because clinical expertise tells us that 120 normally indicates missing information on age. Therefore, although not the actual age of a patient, the value 120 contains information that is valuable to keep.
  • For ethnicity, the invalid inputted values all came in the same format and were handled in the same way, which is better explained by the following example: if category A* was inputted, we assume this meant the correct ethnicity category is A.
@@ 4 Learning causal models
    Construct a knowledge-based structure of the causal relationships between variables present in the data using clinical expertise and existing literature.
@@
    Perform model-averaging on the structures to obtain a model that reflects all algorithms considered.
@@
    Use the CBN for prediction, causal inference and simulation of interventions.
@@ 4.1 Learning the structure of Bayesian Networks
One of the objectives of this study is to investigate possible causes of Sepsis. Structure learning algorithms can be used for this purpose. However, these algorithms cannot generally discover error-free causal structures, mainly when applied to real-world data that tend to violate some of the assumptions the algorithms make about the input data ([23]). Moreover, the following assumptions are required to interpret learnt structures as causal structures that can be converted into CBNs.
  • Causal Markov condition: nodes are independent of every other node, except their descendants, given their parents.
  • Causal faithfulness: there are no independencies in the underlying probability distribution that are not those implied by the DAG.
  • Causal sufficiency: the observed variables include all the common causes of variables in the data.
The above assumptions are often difficult to test in practice and are unlikely to hold for most real-world datasets, including the data we use in this study. However, it is necessary to assume these conditions to perform causal inference.
@@
  • Constraint-based: algorithms that rely on statistical tests of conditional independence to evaluate the hypothesis that variables are independent of each other, and therefore, no edge must exist between them. They also attempt to orientate edges between dependent variables.
@@
For a comprehensive review of structure learning algorithms, please refer to [24], [25] and [26]. The algorithms used in this study are described below by learning class. In the past, constraint-based algorithms were considered more appropriate than score-based algorithms in discovering causal structures. However, recent empirical studies have shown that score-based and hybrid algorithms are often better than constraint-based algorithms for this task ([23], [27]). However, it remains unclear under what circumstances one class of structure learning algorithms might be more appropriate than another, which is why we investigate all three types of algorithms in this study.
@@ 4.1.1 Constraint-based learning
We use the well-established PC-Stable algorithm ([28]), which is an extension of the PC algorithm ([29]) that relaxes the PC’s output sensitivity to the order of the variables as read from data. It starts from a fully connected undirected graph and checks whether each pair of variables is independent or conditionally independent. When variables are independent or conditionally independent, the edge connecting them is removed. When these tests are finished, the algorithm will attempt to orientate the edges preserved in the graph. The algorithm returns a graphical structure containing both directed and undirected edges, reflecting the theoretical limitations of orientating all edges from observational data alone.
@@ 4.1.2 Score-based learning
We use the classic Hill-Climbing (HC) ([31]) algorithm, which represents the simplest type of search. The algorithm begins with an empty graph and proceeds to add, remove or re-direct edges to maximise a given objective function. When the score no longer improves, the search stops and returns the highest-scoring graph. The primary weakness of HC is that it gets stuck at the first local maxima solution it visits.”.
@@ 4.1.3 Hybrid algorithms
In addition to MMHC we also test the H2PC algorithm by [33]. H2PC first constructs the skeleton of the graph, focusing on avoiding false missing edges to learn the local structure around variables. It then performs HC to find the optimal structure.
@@ 4.2 Model averaging
    Rank the edges according to the number of times they appear in the learnt graphs (out of 6).
@@
    Add the edges to the overall graph following the ranking system.
@@
    If an edge has two possible orientations with equal rank, use knowledge to determine its orientation.
@@
    If an edge added creates a cycle, reverse the edge and attempt to add it again.
@@
    If the reversed edge also creates a cycle, delete it.
A decision remains implicit in this averaging method: How many edges will we include in the average graph? We need a cut-off point for the averaging process, which means we need to decide if we will add the edges that appear in at least one of the structures, edges that appear in all of the structures, or anything in between. To make this decision, we will optimize the BIC score, and the averaging process will be parametrized by the number of structures an edge must appear in to be added to the average model. To be clear, the BIC only determines how we filter the averaging process described above, i.e. if edges that appear in 1,2,…, 6 structures will be added to the averaged model. In this study, the averaging process is optimized when we include edges that appear in at least two structures. The BIC score, which we describe in section 4.4, is a statistical measure for model selection that balances model fit with complexity.

### v2

@@ Abstract
Sepsis is a life-threatening and serious global health issue. This study combines knowledge with available hospital data to investigate the potential causes of Sepsis that can be affected by policy decisions. We investigate the underlying causal structure of this problem by combining clinical expertise with score-based, constraint-based, and hybrid structure learning algorithms. A novel approach to model averaging and knowledge-based constraints was implemented to arrive at a consensus structure for causal inference. The structure learning process highlighted the importance of exploring data-driven approaches alongside clinical expertise. This includes discovering unexpected, although reasonable, relationships from a clinical perspective. Hypothetical interventions on Chronic Obstructive Pulmonary Disease, Alcohol dependence, and Diabetes suggest that the presence of any of these risk factors in patients increases the likelihood of Sepsis. This finding, alongside measuring the effect of these risk factors on Sepsis, has potential policy implications. Recognising the importance of prediction in improving health outcomes related to Sepsis, the model is also assessed in its ability to predict Sepsis by evaluating accuracy, sensitivity, and specificity. These three indicators all had results around 70%, and the AUC was 80%, which means the causal structure of the model is reasonably accurate given that the models were trained on data available for commissioning purposes only.
@@ 1 Introduction
Extensive research has been conducted in building predictive models for the early diagnosis of Sepsis. Some examples include [4], [5], [6], [7], [8] and [9]. These models, which include Bayesian Networks and some constructed using supervised learning models, are intended to aid the management of Sepsis in hospitals and prevent severe cases by supporting the early detection of Sepsis. Although these models may help to reduce the incidence and mortality of Sepsis due to their early prediction capabilities, they do not investigate potential areas for intervention to reduce the incidence of Sepsis. Furthermore, all the predictive models reviewed use biomarkers or clinical data for predicting Sepsis. However, these data are not routinely collected by the National Health Service (NHS) in England for service commissioning purposes. This lack of data may be true for commissioners of services in other healthcare systems. Accessing patient records without a clinical or direct care reason is generally restricted. Even if commissioners could access these data, they probably would not have the necessary expertise to interpret it if they do not have a clinical background. Because commissioners play a crucial role in service design, building a model using data they regularly access, like the one developed in this study, is important for exploring new services that can improve patient outcomes.
@@
The methods used for prediction in the public health sector are generally regression-based, as described in [8] and [9]. One problem with regression models is that they do not enable accurate modelling of the often-interdependent relationships of the risk factors for patients developing Sepsis. The relationships between these factors might be highly complex, and it is essential to model them as such.
@@
Regression models do not generally explore possible causal relationships. In contrast, this study examines graphs learnt from data and explores their potential to reveal causal relationships. However, these algorithms rely on assumptions which often do not hold in practice. Here, we attempt to reduce the resulting errors by incorporating expert knowledge and averaging models produced by different algorithms.
@@
Causal Bayesian Networks (CBNs) are represented by a Directed Acyclic Graph (DAG), which provides an intuitive and easy way to read the causal relationships between risk factors and model the impact of interventions available to policymakers. DAGs consist of nodes representing variables and arcs (or edges) assumed to represent causal relationships. The acyclicity condition in DAGs formalises the principle that a variable cannot cause itself. The DAG representation in CBNs provides a clear visualisation of complex clinical relationships and an intuitive framework for identifying causal pathways. The structure learning process described in the previous paragraph determines the structure of these arcs and nodes. One of the relevant features of CBNs is that they enable us to model predictive and diagnostic inference and the effect of hypothetical interventions using the so-called do-calculus framework proposed by [10].
@@
  • Develops a CBN model using data routinely collected by the NHS in England for commissioning purposes.
  • Captures the highly complex interrelations amongst risk factors with a novel method combining expert knowledge and averaging six structure learning algorithms.
  • Assesses the ability of the model to predict the presence of Sepsis.
  • Identifies which risk factors can be influenced by policy decisions and uses algorithms to discover and estimate their causal effects on Sepsis.
Unlike purely predictive methods, the proposed CBN approach reveals how modifiable risk factors can be targeted to reduce the incidence of Sepsis. This focus on actionable insights is critical in policymaking contexts, where commissioners require evidence of which interventions can yield the most significant impact on patient outcomes. In addition, we explore averaging structures across multiple data-driven structures to reduce uncertainty and use expert domain knowledge to judge disagreements between algorithms, thereby mitigating the limitations of purely algorithmic solutions. These limitations originate from the algorithmic reliance on assumptions that may not hold in real-world healthcare settings.
This work bridges a critical gap between technical innovation and practical applicability by introducing a causally grounded, policy-focused approach to Sepsis. The result is a well-suited framework for healthcare commissioners, equipping them with predictive capabilities and causal insights to inform strategic resource allocation and service design. Although built explicitly for Sepsis, we hope the framework used to create this model inspires researchers to apply it to other diseases.
@@ 2.1 Factors Associated with Sepsis
A range of risk factors is identified in the literature as being associated with the presence of Sepsis. The risk factors described below can be found in routinely collected data for service commissioning purposes by the NHS. The discussion did not include risk factors that cannot be found in these data. Table 1 lists these factors and studies identifying them as likely to be linked to Sepsis. Our study will consider all of these factors.
@@
Table 2 presents other variables that we incorporated based on clinical expertise in this study. Clinical expertise was elicited through collaboration with Dr Jonathon Dean, a clinical expert based at the North East London Adult Critical Care Operational Delivery Network and the co-founder of the North East London Critical Care Transfer and Retrieval service. Dr Dean has spent the last five years working in anaesthetics and critical care in North East London, where Sepsis is an important and topical pathology.
@@ 3 Data pre-processing
This study will use the Secondary Uses Service (SUS+) database, which includes data for patients admitted to hospitals (inpatients), patients using accident and emergency services, and patients with outpatient appointments within the NHS in England. This study will look at admissions classified as emergencies; this means admissions to hospitals that were not planned. ICD-10 (International Classification of Diseases, 10th Revision)is the 10th version of the global standard for diagnosis classification ([19]). These codes are used to classify diagnoses in the database and will be crucial to constructing the variables identified by the literature review and expert knowledge. OPCS codes identify the different interventions, operations, and procedures within the NHS in England, and we have used them to create relevant variables in this study. Variables were constructed with clinical guidance, but only non-clinical authors had access to the database. SUS+ was accessed through the National Commissioning Data Repository (NCDR), a pseudonymised patient-level data repository managed by NHS England.
@@ 3.1 Data Extraction
The odbc R package ([20]) was used to connect to the database. This package also allows for the database to be queried using SQL. Each row in the admitted patients table represents an episode during a hospital spell. A spell can be composed of many episodes. The episodes of a spell are likely to contain similar information regarding diagnosis, procedures and patient information. To obtain only one episode representing any given spell, we applied the rule of selecting the most relevant episode during a spell, where the most relevant is determined by the most relevant Healthcare Resource Groups (HRGs) during the spell. In this way, we ensure we are not double counting the same spells while capturing the most important information about that spell. Treatments are grouped into the same HRGs when they are deemed to require similar levels of resources. Therefore, we selected the episode where the highest level of resources was used for treatment.
The data were extracted for the financial year 2019/2020, which ran from the 1st of April 2019 until the 31st of March 2020. Except for the variables representing the Age, number of diagnoses, number of procedures, sex, and ethnicity of a patient, all the variables extracted were in the form of a binary indicator establishing the presence or absence of the risk factor. A sample of 1,000,000 episodes was taken due to server constraints; i.e., the runtime for this study on the server already spans multiple days with this sample size.
@@ 3.2 Missing Data
Structure learning algorithms, including those used in this study, generally require complete datasets. However, the collated data contained missing values, although the amount of missing data is minimal. The proportion of missing values in the variables age, ethnicity, and sex) is less than 0.6%, i.e. fewer than 6,000 rows of our 1,000,000 sample.
A common but often problematic solution to missingness is to delete the rows with missing values. This works well under the very strong assumption that the missing data are Missing Completely At Random (MCAR) and if sufficient sample size remains after deletion. We say that data are MCAR when missingness is unrelated to other variables and to observed and unobserved values of the variable itself. A much more likely assumption to make for missing data in this study is that data are Missing At Random (MAR). Data is MAR when the fact that data is missing is related to other variables in the dataset but not to the variable itself. The data we use in this study is a cross-section of hospital episodes. Therefore, it is more reasonable to assume that the missing data are MAR; i.e., if a missing value is present, it seems plausible that it will be related to other patient information. Because deleting rows with missing data is not a good option for this study, we performed data imputation using the missForest R package ([21], [22])), which is suitable for both types of missingness, MCAR and MAR. This algorithm uses a Random Forest to impute missing values and does not require any parameters to be specified.
A value zero is assigned to a diagnosis or procedure not applicable in a particular case. In our data, we are unable to distinguish whether a missing value represents such a case or whether it has not been collected. Following clinical advice, we have assumed that missing values for diagnosis and procedures represent non-applicability and have assigned a value of zero to missing values for these variables. Where the data records included wrongly inputted values in the variables describing the Age and ethnicity of a person, the following actions were taken:
  • For Age, this was handled by converting the values over 120 to NA (indication of missing data in R) and then used the chosen imputation method. This is because clinical expertise tells us 120 normally indicates missing information on Age. Therefore, although not the actual Age of a patient, the value contains valuable information to retain.
  • For ethnicity, all invalidly inputted values followed the same format and were handled consistently. for example, if category “A*” was inputted, we assume that the correct ethnicity category is “A.”.
@@ 4 Learning causal models
    Construct a knowledge-based structure of the causal relationships between the variables present in the data using clinical expertise and existing literature.
@@
    Perform model-averaging on the learnt structures to obtain a model that reflects all algorithms considered.
@@
    Use the CBN for prediction, causal inference, and simulation of interventions.
@@ 4.1 Learning the structure of Bayesian Networks
One of the objectives of this study is to investigate possible causes of Sepsis. Structure learning algorithms can be used for this purpose. However, these algorithms cannot generally discover error-free causal structures, especially when applied to real-world data, which tend to violate some of the assumptions the algorithms make about the input data [23]. Moreover, the following assumptions are required to interpret learnt structures as causal structures that can be converted into CBNs:
  • Causal Markov condition: nodes are independent of every other node, except their descendants, given their parents. This means that knowing the direct causes of a variable of interest renders any information about its indirect causes or ancestors irrelevant when predicting that variable.
  • Causal faithfulness: there are no independencies in the underlying probability distribution that are not those implied by the DAG. In other words, faithfulness assumes that the statistical dependencies observed in the data align with those dictated by the DAG. When assuming faithfulness, if two variables are independent in the data, this independency should be explainable by the DAG; i.e., the data should not contain any additional independencies beyond those specified by the DAG.
  • Causal sufficiency: the observed variables include all the common causes of the variables in the data. More specifically, if two variables have a hidden common cause, then sufficiency is violated.
The above assumptions are often difficult to test in practice and are unlikely to hold for most real-world datasets, including the data used in this study. However, it is necessary to assume these conditions to perform causal inference.
@@
  • Constraint-based: algorithms that rely on statistical tests of conditional independence to evaluate the hypothesis that variables are independent of each other, and therefore, no edge must exist between them. They also attempt to orient edges between dependent variables.
@@
For a comprehensive review of structure learning algorithms, please refer to [24], [25] and [26]. The algorithms used in this study are described below by learning class. In the past, constraint-based algorithms were considered more appropriate than score-based algorithms for discovering causal structures. However, recent empirical studies have shown that score-based and hybrid algorithms are often better than constraint-based algorithms for this task ([23], [27]). However, it remains unclear under what circumstances one class of structure learning algorithms might be more appropriate than another, which is why we investigate all three types of algorithms in this study.
@@ 4.1.1 Constraint-based learning
We use the well-established PC-Stable algorithm ([28]), which is an extension of the PC algorithm ([29]) that relaxes the PC’s output sensitivity to the order of the variables as read from data. It starts from a fully connected undirected graph and checks whether each pair of variables is independent or conditionally independent. When variables are independent or conditionally independent, the edge connecting them is removed. When these tests are finished, the algorithm attempts to orient the edges that remain in the graph. Ultimately, the algorithm returns a graphical structure containing both directed and undirected edges, reflecting the theoretical limitations of orientating all edges from observational data alone.
@@ 4.1.2 Score-based learning
We use the classic Hill-Climbing (HC) ([31]) algorithm, which represents the simplest type of search. The algorithm begins with an empty graph and proceeds to add, remove or re-direct edges to maximise a given objective function. When the score no longer improves, the search stops and returns the highest-scoring graph. The primary weakness of HC is that it gets stuck at the first local maxima solution it visits.
@@ 4.1.3 Hybrid algorithms
In addition to MMHC, we test the H2PC algorithm by [33]. H2PC first constructs the skeleton of the graph, focusing on avoiding false missing edges to learn the local structure around variables. It then performs HC to find the optimal structure.
@@ 4.2 Model averaging
    Learn the structures: Generate $K$ different structures. In our study, $K=6$.
@@
    Initialise edge list: Collect all edges that appear in the $K$ structures and compile them into a list.
@@
    Count edge frequencies: Record the number of times, $J$, an edge appears across all structures.
@@
    Sort edges by frequency: Rank all edges in descending order based on how frequently they appear across the six graphs.
@@
    Add edges iteratively: Start from the most frequent edges and proceed down the ranking, and add edges to the overall graph.
  6. 6.
    Handle edge orientations: If an edge has two possible orientations with equal rank, use domain knowledge to determine its orientation.
  7. 7.
    Avoid cycles: If adding an edge creates a cycle, reverse the edge and attempt to add it again. If the reversed edge also creates a cycle, remove the edge entirely.
  8. 8.
    Using the edge counts $J$, we create new sets of structures $L$, where each $L$ contains all edges that appear at least $J$ times; e.g., $L=2$ represents the network that includes all edges with $J\geq 2$.
  9. 9.
    Determine optimal averaging: Get the BIC (defined in section 4.4) score for each structure in $L$. In this study, the optimal averaged structure was $L=2$.